# EnvGen: Generating and Adapting Environments via LLMs for Training Embodied Agents

### v1

@@ 4.3 Adaptation of Training Environments Helps the Agent Improve Weaker Skills
Fig. 7 shows how the LLM adaptively generates new training environments based on the intermediate performance of our PPO-based RL agent. In the intermediate performance plots, we compare the baseline agent trained only in Crafter and our RL agent trained in Crafter. In the cycle 2, given the feedback that the current RL agent is not good at collecting coal, the LLM could generate an environment to help the agent focus on the skill, improving the agent performance in ‘collect coal’. Likewise, in the cycle 3, given the feedback that the agent is weak at making stone pickaxe, the LLM could generate an environment to help the agent more easily craft the stone pickaxe, helping the agent improving the score in ‘make stone pickaxe’. Powered by the adaptive LLM environment generation of EnvGen, our agent learns to unlock these two achievements significantly faster than the baseline agent.
| PPO | - | 25M | 25.9 13.2 | 4.1 1.8 |
| PPO + EnvGen (Ours) | 5M | 20M | 37.7 7.50 | 5.5 0.9 |
4.4 Evaluation on Heist Environment
EnvGen can generalize to Heist.
We also evaluate the effectiveness of EnvGen framework with another game environment – Heist, a maze navigation game described in Sec. 3.1. We compare the PPO-based agent trained with and without EnvGen (i.e., Heist environments). Table 2 shows that training an agent with Heist environments is effective in improving performance by increasing average scores (25.9% 37.7%) and rewards (4.1% 5.5%), while also stabilizing training by reducing the score variance (i.e., standard deviation goes down 13.2% 7.5%).
@@ 4.5 Additional Analysis and Ablation Studies
In the following, we show comprehensive ablation studies of EnvGen method: EnvGen vs. longer training in the original environment, dynamically updating LLM environments (i.e., using adaptive environments) vs. using a fixed LLM environment, different LLMs for generating environments, frequency of environment updates, the number of LLM environments, and the ratio of training steps in the LLM environments to the original environment. Unless otherwise noted, we use the PPO-based agent (Moon et al., 2023) (described in Sec. 3.2) on the Crafter (Hafner, 2022) benchmark (described in Sec. 3.1) with 0.96M steps in Crafter and average results for 30 runs (10 different seeds, 3 different initial environments).
@@
| Crafter environments during training | Score (%) | Reward |
| Fixed | 29.9 0.9 | 12.6 0.8 |
| Updated based on RL agent performance (default) | 32.2 0.6 | 12.6 0.6 |
Adaptive environments: Fixed. vs. Updated based on RL agent performance.
Table 4 shows that using LLM environments that are adaptively updated based on intermediate agent performance to improve weaker skills results in overall higher scoring agents than just using the initial LLM environments for the whole training (32.2% vs. 29.9%). These results indicate the effectiveness of the agent feedback and environment updating (step 4 described in Sec. 2).
@@
| Environment Update Frequency | # Training cycles | Score (%) | Reward |
| Every 0.012M steps | 40 cycles | 30.8 0.7 | 12.8 0.6 |
| Every 0.06M steps | 8 cycles | 32.1 0.5 | 12.7 0.8 |
| Every 0.12M steps (default) | 4 cycles | 32.2 0.6 | 12.6 0.6 |
Frequency of LLM feedback / environment updates.
Table 6 shows that updating the LLM environments at every 0.12M steps results in the best agent performance. While increasing the cycles of environment feedback beyond 4 does not improve further, we find that updating environments with feedback always helps improve the RL agent’s performance compared to training only with the original Crafter environment (26.4%) or the fixed LLM environment (29.9%).
@@
5 Related Works
LLMs as open-world game agents.
As LLMs have shown rapid progress in various domains (Brown et al., 2020; OpenAI, 2023a; Touvron et al., 2023a; b; Chowdhery et al., 2023; Anil et al., 2023; Gemini Team, 2023), recent works study using LLMs to create action plans (i.e., a list of subgoals or skills to target) for embodied agents in open-world games like Minecraft and Crafter (Hafner, 2022). Most of these methods require calling LLMs frequently (e.g., at every step) for planning the next steps (Yuan et al., 2023; Wang et al., 2023c; Wu et al., 2023; Wang et al., 2023a; d; Zhao et al., 2023). Other methods, such as Li et al. (2024); Kwon et al. (2023); Ma et al. (2023); Du et al. (2023), have used LLMs to create/adjust rewards to train agents. Although these works show initial promising results leveraging the world knowledge of LLMs to tackle long-horizon tasks, iteratively calling LLMs throughout episodes is prohibitively slow and expensive (e.g., running a single episode in the Crafter environment with SPRING (Wu et al., 2023) costs around $270 USD as they have 2.7K LLM calls on average). EnvGen framework proposes an alternative: calling LLMs only a few times (e.g., 4) throughout the learning process to create training environments that focus on helping the RL agent progressively improve the skills that the agent is weak at.
Reward designs in reinforcement learning.
Finding good action trajectories is critical in reinforcement learning (RL) (Sutton & Barto, 2018). While classic random exploration algorithms such as epsilon-greedy (Watkins, 1989) work well in simple settings such as multi-armed bandit, it is not the case for hard exploration problems where the environment gives very sparse rewards (Weng, 2020). A line of work studies how to augment the original (extrinsic) rewards from the environment with intrinsic rewards that encourage exploration (Bellemare et al., 2016; Burda et al., 2018). While such intrinsic rewards can help RL agents discover novel states and improve their knowledge about the environment, it often requires long pretraining and does not guarantee that the intrinsic reward can help the target task. Another recent line of work studies using LLMs to adjust reward functions to help RL agents progressively learn certain tasks (Li et al., 2024; Kwon et al., 2023; Ma et al., 2023; Du et al., 2023). Instead of designing new rewards, in EnvGen, an LLM adaptively generates training environments that can help the RL agent learn multiple skills it is weak at with fewer training steps than in the original environment; reward design could be complementary to our method.
Deep learning-based game/simulator content generation.
Procedural content generation (PCG) for games is about the automatic generation of levels, landscapes, items, rules, quests, or other types of game contents (Shaker et al., 2016). While traditional PCG methods are based on search/solver/rule/grammar-based methods, recent works study applying deep learning methods such as GAN (Goodfellow et al., 2014) for PCG (Liu et al., 2021; Kumaran et al., 2020; Schubert et al., 2022). Several works have recently explored using LLMs to generate game content such as difficulty levels (Sudhakaran et al., 2023; Todd et al., 2023) and scenes/environments (Kumaran et al., 2023; Wang et al., 2023b; Afshar & Li, 2024). While these works study how to help developers create extra/new game content in simulators, generated environments may not be good for teaching agents how to play/be better in the original environment. Our work studies using LLMs to adaptively generate training environments that teach multiple useful skills to produce better-performing embodied RL agents. Beyond game content generation, several recent works investigate visually augmenting vision-and-language navigation (VLN) simulators (e.g., rendering the environments with different styles) using image generation models (Li et al., 2022b; Wang et al., 2023e; Li & Bansal, 2023). Such works could complement our LLM-based environment generation (e.g., rendering our LLM-generated environments with diverse colors and textures).
6 Conclusion
We propose EnvGen, a novel framework to improve embodied agent performance by utilizing the world knowledge of LLMs to adaptively generate customized training environments that progressively teach agents different skills more effectively. In EnvGen, we give an LLM a prompt describing a game/simulator and ask the LLM to generate the configurations to create new environments that can teach different skills. Next, we train the agents in the LLM-generated environments and give feedback to the LLM by testing the agent performance in the original environments, and then ask the LLM to update the environments to teach agents skills they are weaker at. We experiment with two different games, Crafter and Heist, and find that our EnvGen framework effectively can increase agent performance. We also demonstrate that training in LLM-generated environments can be more effective than just training longer in the original environments when learning long-horizon tasks. Moreover, we show that a lightweight model ( 5M parameters) trained with LLM-generated environments can even outperform an LLM agent, with significantly fewer LLM calls. We also show comprehensive analyses of our results and ablation studies validating the design choices of EnvGen framework. We hope our work can guide future works in leveraging LLMs for training embodied agents.
Acknowledgments
We thank Elias Stengel-Eskin for the thoughtful discussion. This work was supported by DARPA ECOLE Program No. HR00112390060, NSF-AI Engage Institute DRL-2112635, DARPA Machine Commonsense (MCS) Grant N66001-19-2-4031, ARO Award W911NF2110220, ONR Grant N00014-23-1-2356, and a Bloomberg Data Science Ph.D. Fellowship. The views contained in this article are those of the authors and not of the funding agency.
Appendix
In this appendix, we present additional experiment results with another RL agent (Appendix A), RL agent implementation details (Appendix B), additional LLM details (Appendix C), and limitations (Appendix D).
Appendix A Additional Experiment Results with Another RL Agent
@@
Appendix B RL Agent Implementation Details
@@
Appendix C Additional LLM Details
@@ Appendix D Limitations
EnvGen also requires that the environment simulators can (or be easily edited to) accept configurations in standard formats (e.g., JSON, CSV, YAML, TOML etc.), and the LLM can correctly generate configurations in such formats. Note that such text configuration formats are widely used for managing game simulators. We empirically find that environment configurations generated by GPT-4 can be easily parsed without error.

### v2

@@ 4.3 Adaptation of Training Environments Helps the Agent Improve Weaker Skills
Fig. 6 shows how the LLM adaptively generates new training environments based on the intermediate performance of our PPO-based RL agent. In the intermediate performance plots, we compare the baseline agent trained only in Crafter and our RL agent trained in Crafter. In the cycle 2, given the feedback that the current RL agent is not good at collecting coal, the LLM generates an environment to help the agent focus on it, improving the agent’s performance for the skill. Likewise, in the cycle 3, given the feedback that the agent is weak at making stone pickaxes, the LLM generates an environment to help the agent more easily craft the stone pickaxe, helping the agent improve it’s score for the skill. Powered by the adaptive LLM environment generation of EnvGen, our agent learns to unlock these two achievements significantly faster than the baseline agent.
@@ 4.4 Additional Analysis and Ablation Studies
In the following, we show comprehensive design analysis and ablation studies of EnvGen method: dynamically updating LLM environments (i.e., using adaptive environments) vs. curriculum learning methods, and different frequencies of environment updates. In Appendix C, we show comprehensive analysis and ablation studies of EnvGen method: EnvGen vs. longer training in the original environment, different LLMs for generating environments, the number of LLM environments, and the ratio of training steps in the LLM vs. original environments. We also include experiments on the Heist environment (see Sec. 3.1) in Sec. C.2.
| Training Curriculum | Score (%) | Reward |
| Fixed (no curriculum) | 29.9 0.9 | 12.6 0.8 |
| Easy-to-Hard | 26.8 1.5 | 12.7 0.7 |
| Adversarial | 26.8 0.8 | 12.2 0.7 |
| Adaptive+Dynamic Environments (EnvGen) | 32.2 0.6 | 12.6 0.6 |
Different environment curricula: fixed, easy-to-hard, adversarial vs. adaptive. Table 2 shows that using LLM environments that are adaptively updated based on intermediate agent performance to improve weaker skills (last row) results in overall higher scoring agents than just using the initial LLM environments for the whole training (32.2% vs. 29.9%). These results indicate the effectiveness of the agent feedback and environment updating (step 4 described in Sec. 2).
Table 2 also compares an agent trained via EnvGen to the same agent trained with curriculum learning approaches such as an easy-to-hard curriculum, similar to Ammanabrolu et al. (2022) (i.e., pre-defined training environment order based on environment difficulty) and adversarial curriculum, similar to Parker-Holder et al. (2022) (i.e., updating to training environments that agent does worse in) in the Crafter environment. Detailed setups of both baseline approaches are in the appendix. The agent trained with EnvGen is able to achieve much higher performance (32.2% vs. 26.8% for both curricula) indicating the effectiveness EnvGen’s approach of adaptively generating training environments to improve agent weak skills. The result indicates that creating more difficult environments does not necessarily help the agent learn new skills over time.
| Environment Update Frequency | # Training cycles | Score (%) | Reward |
| Every 0.012M steps | 40 cycles | 30.8 0.7 | 12.8 0.6 |
| Every 0.06M steps | 8 cycles | 32.1 0.5 | 12.7 0.8 |
| Every 0.12M steps (default) | 4 cycles | 32.2 0.6 | 12.6 0.6 |
Frequency of LLM feedback / environment updates.
Table 3 shows that updating the LLM environments at every 0.12M steps results in the best agent performance. While increasing the cycles of environment feedback beyond 4 does not improve further, we find that updating environments with feedback always helps improve the RL agent’s performance compared to training only with the original Crafter environment in Table 1 (26.4%) or the fixed LLM environment in Table 2 (29.9%).
5 Related Works
LLMs as open-world game agents. Recent works study using LLMs to create action plans (i.e., a list of subgoals or skills to target) for embodied agents in open-world games like Minecraft and Crafter (Hafner, 2022). Most of these methods require calling LLMs frequently (e.g., at every step) for planning the next steps (Yuan et al., 2023; Wang et al., 2023c; Wu et al., 2023; Wang et al., 2023a; d; Zhao et al., 2023). Other methods, such as Li et al. (2024); Kwon et al. (2023); Ma et al. (2023); Du et al. (2023), have used LLMs to create/adjust rewards to train agents. Although these works show initial promising results leveraging the world knowledge of LLMs to tackle long-horizon tasks, iteratively calling LLMs throughout episodes is prohibitively slow and expensive (e.g., running a single episode in the Crafter environment with SPRING (Wu et al., 2023) costs around $270 USD as they have 2.7K LLM calls on average). EnvGen only calls LLMs a few times (e.g., 4 in total) to create training environments that focus on helping the RL agent progressively improve its weaker skills.
Deep learning-based game/simulator content generation. Procedural content generation (PCG) for games is about the automatic generation of levels, landscapes, items, rules, quests, or other types of game contents (Shaker et al., 2016). While traditional PCG methods are based on search/solver/rule/grammar-based methods, recent works apply deep learning methods such as GAN (Goodfellow et al., 2014) for PCG (Liu et al., 2021; Kumaran et al., 2020; Schubert et al., 2022). Several works have explored using LLMs to generate game content such as difficulty levels (Sudhakaran et al., 2023; Todd et al., 2023) and scenes/environments (Kumaran et al., 2023; Wang et al., 2023b; Afshar & Li, 2024). While these works aim to help developers create new game content, we aim to improve RL agent performance in the original environment. A line of work proposes unsupervised environment design (UED) that manipulates the difficulty level of environments to be more challenging to RL agents (Dennis et al., 2020; Jiang et al., 2021; Parker-Holder et al., 2022). While these works use a learned environment manipulator or evolutionary algorithms to maximize the ‘regret’ (the difference between the expected return of the current and optimal policies) in simple games such as MiniGrid (Chevalier-Boisvert et al., 2023), we use the world knowledge of LLMs to generate and adapt training environments that can improve weaker skills based on comprehensive skill-specific feedback from RL agents in open-world games with many challenging long-horizon tasks. To help agents generalize to unseen tasks in a text-based dialogue game, Ammanabrolu et al. (2022) augment new tasks with LMs and use a manually designed, fixed curriculum. Unlike this work, we adaptively generate training environments using LLMs’ world knowledge and automatically learning a dynamic curriculum based on the RL agent’s feedback, so as to improve the agent’s weaker skills in open-world games with visual inputs. Beyond game content generation, several works visually augment vision-and-language navigation (VLN) simulators (e.g., rendering the environments with different styles) using image generation models (Li et al., 2022b; Wang et al., 2023e; Li & Bansal, 2023). Such works could complement our LLM environments (e.g., augmenting our environments with diverse colors and textures).
6 Conclusion
We propose EnvGen, a novel framework to improve embodied RL agent performance by utilizing the world knowledge of LLMs to adaptively generate training environments. In EnvGen, we give an LLM a prompt describing a game/simulator and ask the LLM to generate the configurations to create new environments that can teach different skills. Next, we train an agent in the LLM-generated environments, give feedback to the LLM by testing the agent in the original environments, and then ask the LLM to update the environments to teach agents skills they are weaker at. In two challenging games, Crafter and Heist, we show that EnvGen increases agent performance significantly, and training with LLM-generated environments is more effective than training longer in the original environments. We also show that using an LLM to adapt environments dynamically outperforms curriculum learning approaches and how the LLM adapts training environments to help improve RL agents’ weaker skills over time. Moreover, a lightweight model ( 5M parameters) trained with LLM-generated environments even outperforms an LLM agent with significantly fewer LLM calls. We hope our work can guide future works in leveraging LLMs for embodied agents.
Acknowledgments
We thank Elias Stengel-Eskin and the reviewers for the thoughtful discussion and feedback. This work was supported by DARPA ECOLE Program No. HR00112390060, NSF-AI Engage Institute DRL-2112635, DARPA Machine Commonsense (MCS) Grant N66001-19-2-4031, ARO Award W911NF2110220, ONR Grant N00014-23-1-2356, and a Bloomberg Data Science Ph.D. Fellowship. The views contained in this article are those of the authors and not of the funding agency.
Appendix
In this appendix, we present additional related work (Appendix A), additional game environment details (Appendix B), additional experiment results (Appendix C), curriculum learning baseline method details (Appendix D), RL agent implementation details (Appendix E), additional LLM details (Appendix F), and limitations (Appendix G).
Appendix A Additional Related Works
Reward designs in reinforcement learning. Finding good action trajectories is critical in reinforcement learning (RL) (Sutton & Barto, 2018). While classic random exploration algorithms such as epsilon-greedy (Watkins, 1989) work well in simple settings such as multi-armed bandit, it is not the case for hard exploration problems where the environment gives very sparse rewards (Weng, 2020). A line of work studies how to augment the original (extrinsic) rewards from the environment with intrinsic rewards that encourage exploration (Bellemare et al., 2016; Burda et al., 2018). While such intrinsic rewards can help RL agents discover novel states and improve their knowledge about the environment, it often requires long pretraining and does not guarantee that the intrinsic reward can help the target task. Another recent line of work studies using LLMs to adjust reward functions to help RL agents progressively learn certain tasks (Li et al., 2024; Kwon et al., 2023; Ma et al., 2023; Du et al., 2023). Instead of designing new rewards, in EnvGen, an LLM adaptively generates training environments that can help the RL agent learn multiple skills it is weak at with fewer training steps than in the original environment; reward design could be complementary to our method.
Appendix B Additional Game Environment Details
Heist Environment.
Appendix C Additional Experiment Results
C.1 Design Choices, Ablations, and Other Agent Architectures
In the following, we show comprehensive design choice and ablation studies of EnvGen method: EnvGen vs. longer training in the original environment, different LLMs for generating environments, the number of LLM environments, and the ratio of training steps in the LLM environments to the original environment. Unless otherwise noted, we use the PPO-based agent (Moon et al., 2023) (described in Sec. 3.2) on the Crafter (Hafner, 2022) benchmark (described in Sec. 3.1) with 0.96M steps in Crafter and average results for 30 runs (10 different seeds, 3 different initial environments).
@@ C.1 Design Choices, Ablations, and Other Agent Architectures
Can simulators always understand LLM-generated environment configurations?
We tested and analyzed the generated environments used in paper experiments (109 total) and found that the LLM (GPT-4-Turbo) did not generate any environments beyond what the Crafter or Heist simulators could handle. While we do not find any such case, even if the LLM generates an invalid setup, we constrain all LLM-generated settings in post-processing to be within simulator capabilities to ensure no accidental errors in the simulator or environment generation.
Environment parameter naming: obscure vs. original.
To determine whether or not the LLM in EnvGen is leveraging world knowledge when generating environments we conduct an analysis experiment. We replace environment parameter names from the original ones with obscure names (see Fig. 9), in order to remove the use of prior knowledge/expectation of how each parameter is useful to which skills. We find that the performance decreases, from 32.2 0.6 to 28.5 0.6, indicating that the world knowledge/prior knowledge of the LLM is beneficial in helping the LLM generate more suitable environments.
@@
C.2 Evaluation on Heist Environment
| PPO | - | 25M | 25.9 13.2 | 4.1 1.8 |
| PPO + EnvGen (Ours) | 5M | 20M | 37.7 7.50 | 5.5 0.9 |
EnvGen can generalize to Heist.
We also evaluate the effectiveness of EnvGen framework with Heist. We compare the PPO-based agent trained with and without EnvGen (i.e., Heist environments). Table 8 shows that training an agent with Heist environments is effective in improving performance by increasing average scores (25.9% 37.7%) and rewards (4.1% 5.5%), while also stabilizing training by reducing the score variance (i.e., standard deviation goes down 13.2% 7.5%).
Before:
Here is a list of parameters you can control when making an environment:
target_biome: grassland | mountain | beaches | natural
coal_rarity: very common | common | rare
iron_rarity: very common | common | rare
diamond_rarity: very common | common | rare
tree_rarity: very common | common | rare
Here is a list of items the agent can start with:
sapling: 0-9
wood: 0-9
stone: 0-9
coal: 0-9
iron: 0-9
diamond: 0-9
wood_pickaxe: 0-1
stone_pickaxe: 0-1
iron_pickaxe: 0-1
wood_sword: 0-1
stone_sword: 0-1
iron_sword: 0-1
After:
Here is a list of parameters you can control when making an environment:
parameter1: optionA | optionB | optionC | optionD
parameter2: optionE | optionF | optionG
parameter3: optionE | optionF | optionG
parameter4: optionE | optionF | optionG
parameter5: optionE | optionF | optionG
Here is a list of items the agent can start with:
item1: 0-9
item2: 0-9
item3: 0-9
item4: 0-9
item5: 0-9
item6: 0-9
item7: 0-1
item8: 0-1
item9: 0-1
item10: 0-1
item11: 0-1
item12: 0-1
Appendix D Curriculum Learning Baseline Details
In the following, we describe the two baseline implementation details described in Table 2: easy-to-hard and adversarial curricula.
Easy-to-hard curriculum.
Similar to Ammanabrolu et al. (2022), we create an easy-to-hard curriculum. An easy-to-hard curriculum has a pre-defined order of training environments. The agent first trains in “easy" environments and then by the end of the training process will be training in the “hard" environments. To do this, we first ask the LLM to generate a set of 16 random environments and train a validation agent (an agent only for the purpose of testing an environment difficulty; not used during final agent training) on each environment. Then the performance of the validation agent is measured and the environments are sorted from easiest to hardest (i.e., from environments that resulted in higher agent scores to lower agent scores). Then we train an agent on the top four easier environments first and for every 0.24M steps we replace the training environments with the next four environments in the sorted set (i.e., easy-to-hard curriculum).
Adversarial curriculum.
Similar to Parker-Holder et al. (2022), we create an adversarial curriculum. The adversarial curriculum approach involves updating the agent’s training environments to ones where it has struggled. To do this, we first generate a set of 16 random environments and train a validation agent on each environment. Then, we measure and sort the environments by difficulty (i.e., sorted from lowest to highest based on the validation agent’s score). We take the top four hardest environments and train the final agent on these. Then, generate a new set of environments and test the current agent on this set, again sorting by difficulty. Then again, we take the top four hardest environments and resume training on these. This process then repeats four times (every 0.24M steps) creating an adversarial curriculum.
Appendix E RL Agent Implementation Details
@@
Appendix F Additional LLM Details
@@ Appendix G Limitations
EnvGen also requires that the environment simulators can (or be easily edited to) accept configurations in standard formats (e.g., JSON, CSV, YAML, TOML etc.), and the LLM can correctly generate configurations in such formats. Note that such text configuration formats are widely used for managing game simulators. In addition, many games have open-source community-driven efforts that provide high-level configuration documentation and settings, such as Minecraft wrappers (Guss et al., 2019; Fan et al., 2022) and Starcraft wrappers (Vinyals et al., 2017; DI-star Contributors, 2021).